OpenStudy (anonymous):

Write the repeating decimal 0.3181818 as a simplified fraction. I need to show work but I'm not sure how

OpenStudy (campbell_st):

let the decimal = x 1000x = 318.181818... (1) 10x = 3.18181818.... (2) now use equation (1) - Equation 92) 990x = 315 x = 315/990 simplify x = 7/22

OpenStudy (anonymous):

oops, i did'nt see it as a repeating 1818181818188181818188181818181

OpenStudy (campbell_st):

the subtraction process eliminates the recurring decimal